Check whether the function f given by



f(x)=1/(2x-4)^2 ∀x∈]-2,2[



Is continuous in the interaval ]-2,2] but not bounded in the interval (-2,2)

1
Expert's answer
2022-03-18T05:33:14-0400

The function is continuous since it is rational function with the denominator non-zero on (-2,2). It is not bounded since "f(2-\frac{1}{2n})=n^2\to +\infty, n\to\infty"